# LangSplatV2: High-dimensional 3D Language Gaussian Splatting with 450+ FPS

### v1

@@ Appendix B Discussion
LangSplatV2 proposes to model the high-dimensional language features of each Gaussian point as a sparse code within a global dictionary. While similar concepts, such as quantizing Gaussian point attributes (e.g., SH coefficients fan2023lightgaussian , opacity niedermayr2024compressed , and semantic features shi2024language ) into a codebook, have been explored in the domains of compression niedermayr2024compressed and language embedding shi2024language , our method is fundamentally different.
In the compression domain, existing approaches niedermayr2024compressed ; fan2023lightgaussian first learn unique attributes for each Gaussian point, which are then quantized into a codebook using a quantizer. Extending this concept to our scenario would require learning 512-dimensional semantic features for each Gaussian point before quantization. This process significantly increases memory consumption and training time, often exceeding the capacity of GPUs like the NVIDIA 3090 or 4090, leading to out-of-memory (OOM) errors. Moreover, even with advanced GPUs that can handle the training, the rendering process would still involve managing 512-dimensional features. As shown in Figure 1 of the main paper, this dramatically reduces rendering speed compared to LangSplatV2.
@@
Furthermore, due to the reliance on an MLP decoder, LEGaussians still operate in a lower-dimensional feature space, which limits the expressiveness of its learned representations. Additionally, its two-stage 2D codebook approach fails to effectively incorporate 3D priors, resulting in a suboptimal learned field. In contrast, LangSplatV2 directly models high-dimensional features within a global 3D codebook, effectively capturing the underlying 3D structures. As evidenced by the results in Table 7 of the main paper, our approach outperforms LEGaussians across multiple benchmarks while being significantly faster, demonstrating superior scene reconstruction quality and more accurate language-grounded 3D representations.
@@
Appendix C More Ablation Study
@@
Method Training Time (h) Training GPU Memory (GB)
LangSplat qin2023langsplat 1.0 6.2
LEGaussian shi2023language 1.3 11
LangSplatV2 3.0 21.2
LangSplat qin2023langsplat (512 Dim) 45.0 47.4
Figure 6: More qualitative comparisons of open-vocabulary 3D object localization on the LERF dataset. The red points are the model predictions and the black dashed bounding boxes denote the annotations. We observe that LangSplatV2 generates better results than LangSplat.
Training Cost. Table 11 shows the comparison of training cost on the LERF kerr2023lerf dataset with one A100 GPU. Compared with LEGaussians shi2023language and LangSplat qin2023langsplat , LangSplatV2 comes with increased training cost due to the need to construct high-dimensional semantic fields during training. Although our LangSplatV2 incurs higher training costs compared to LangSplat and LEGaussians, our primary focus in this paper is on improving the model’s test-time performance for deployment. As demonstrated in the paper, LangSplatV2 achieves significantly higher inference speed—47$\times$ faster than LangSplat and 14$\times$ faster than LEGaussians on the LERF dataset—under comparable memory consumption, while also yielding higher segmentation accuracy. Furthermore, compared with naive LangSplat (directly trained with 512 dimensions), LangSplatV2 is more efficient with the proposed global codebook and sparse coefficient field.
Appendix D More Visualization Results
Open-vocabulary 3D Object Localization. We visualize more examples on the LERF kerr2023lerf and Mip-NeRF360 kerr2023lerf datasets for open-vocabulary 3D object localization in Fig 6.
Open-vocabulary 3D Semantic Segmentation. We visualize more examples on the LERF kerr2023lerf , Mip-NeRF360 JTBarron2022mip and 3D-OVS liu2023weakly datasets for open-vocabulary 3D semantic segmentation in Fig 7.

### v2

@@ Appendix B Discussion
LangSplatV2 proposes to model the high-dimensional language features of each Gaussian point as a sparse code within a global dictionary. While similar concepts, such as quantizing Gaussian point attributes (e.g., SH coefficients fan2024lightgaussian , opacity niedermayr2024compressed , and semantic features shi2024language ) into a codebook, have been explored in the domains of compression niedermayr2024compressed and language embedding shi2024language , our method is fundamentally different.
In the compression domain, existing approaches niedermayr2024compressed ; fan2024lightgaussian first learn unique attributes for each Gaussian point, which are then quantized into a codebook using a quantizer. Extending this concept to our scenario would require learning 512-dimensional semantic features for each Gaussian point before quantization. This process significantly increases memory consumption and training time, often exceeding the capacity of GPUs like the NVIDIA 3090 or 4090, leading to out-of-memory (OOM) errors. Moreover, even with advanced GPUs that can handle the training, the rendering process would still involve managing 512-dimensional features. As shown in Figure 1, this dramatically reduces rendering speed compared to LangSplatV2.
@@
Figure 6: More qualitative comparisons of open-vocabulary 3D object localization on the LERF and Mip-NeRF360 datasets. The red points are the model predictions and the black dashed bounding boxes denote the annotations. We observe that LangSplatV2 generates better results than LangSplat.
Furthermore, due to the reliance on an MLP decoder, LEGaussians still operate in a lower-dimensional feature space, which limits the expressiveness of its learned representations. Additionally, its two-stage 2D codebook approach fails to effectively incorporate 3D priors, resulting in a suboptimal learned field. In contrast, LangSplatV2 directly models high-dimensional features within a global 3D codebook, effectively capturing the underlying 3D structures. As evidenced by the results in Table 8, our approach outperforms LEGaussians across multiple benchmarks while being significantly faster, demonstrating superior scene reconstruction quality and more accurate language-grounded 3D representations.
@@
Appendix C Evaluation Details
After calculating the similarity between the rendered semantic features and the CLIP features of the query text, we obtained relevance maps $M$ at different semantic levels. Then, we use different strategies for different datasets to obtain the corresponding query results.
LERF. For object localization, we first apply average pooling to smooth $M$ and get $Avg(M)$. Then we select the point with the highest relevancy in $Avg(M)$ as the predicted point. Additionally, we choose the semantic level corresponding to the highest relevancy value among the three levels as the selected level. For semantic segmentation, after getting $Avg(M)$, we select the level in the same way as object localization. We then normalize $Avg(M)$ and use a threshold to obtain a binary mask as the final prediction mask.
3D-OVS. The post-processing pipeline for semantic segmentation is similar to that used for the LERF dataset. Differently, we obtain the smoothed relevancy map $Avg^{\prime}(M)$ as $Avg^{\prime}(M)=0.5\times(M+Avg(M))$, and the level selection is based on the average relevancy within the predicted mask. Additionally, when evaluating query accuracy, we use two coarse-grained semantic levels (whole and part) owing to the simplicity of the scenes in this dataset.
Mip-NeRF360. Similar to the post-processing for the 3D-OVS dataset, we obtain the smoothed relevancy map $Avg^{\prime}(M)$ and select the semantic level based on the average relevancy score. When evaluating query accuracy, we use all three semantic levels, as the complex scenes in this dataset demand finer-grained language feature rendering.
Appendix D More Ablation Study
@@
Appendix E More Visualization Results
Open-vocabulary 3D Object Localization. We visualize more examples on the LERF kerr2023lerf and Mip-NeRF360 kerr2023lerf datasets for open-vocabulary 3D object localization in Figure 6.
Open-vocabulary 3D Semantic Segmentation. We visualize more examples on the LERF kerr2023lerf , Mip-NeRF360 JTBarron2022mip and 3D-OVS liu2023weakly datasets for open-vocabulary 3D semantic segmentation in Figure 7.